OpenStudy (keana):

Line AB contains points A (1, 2) and B (−2, 6). The slope of line AB is a. zero b. undefined c. positive d. negative

OpenStudy (gorv):

(6-2)/(-2-1)=4/-3=-4/3 negative

OpenStudy (mazefordays090):

negative

OpenStudy (anonymous):

d.negative; (6-2)/(-2-1)= 4/(-3)= -4/3.

OpenStudy (amistre64):

It is better to guide the user to an answer rather than doing all the work yourself. We discourage giving out answers without letting the asker at least take some part in the analysis - the aim of this site is to teach. Please refer to our code of conduct at http://openstudy.com/code-of-conduct for more information.

OpenStudy (keana):

Ohhh! Okay, now I get it! I wasn't sure what formula to use. Thanks everyone!